You have credit card debt of $30,000 that has an APR (monthly compounding) of 16%. Each month you pay the minimum monthly payment only. You are required to pay only the outstanding interest. You have received an offer in the mail for an otherwise identical credit card with an APR of 11%. After considering all your alternatives, you decide to switch cards, roll over the outstanding balance on the old card into the new card, and borrow additional money as well. How much can you borrow today on the new card without changing the minimum monthly payment you will be required to pay? 

Note: Make sure all calculations are held to at least five (5) decimal places.
Answer #1

Monthly Payment on old card = 30,000(0.16/12) = $400

On new card,

Calculating Maximum loan amount on same monthly payment,

Loan Amount = 400/(0.11/12) = $43,636.36

Additional Debt = 43,636.36 - 30,000

Additional Debt = $13,636.36
Answer #2

30000 x 0.16/12 = Debt on new card x 0.11/12
==> Debt on new card = 43636.36364

Additonal credit = 43636.36364 - 30000 = 13636.36364


answered by: K. Lee